# Redefining Superalignment: From Weak-to-Strong Alignment to Human-AI Co-Alignment for Sustainable Symbiotic Society

### v3

@@ 4 Intrinsic Proactive Superalignment
Externally supervised alignment uses mathematical optimization to derive action policies that minimize loss functions, where the supervision signals provided by humans or AI represent the ceiling of alignment. However, AI behavior remains unexplainable and does not reflect genuine understanding of the constraints or supervisory signals. This leads the system to make many unexpected mistakes, and these potential dangers are precisely the kind of catastrophic risks that cannot be covered by external supervisory signals. Thus, in addition to passive external supervision and constraints, we also need to explore internal mechanisms alignment to proactively develop and shape AI, enabling it to spontaneous align with human ethical values beyond mere external supervision.
@@ 4.2 Beneficial Meaningful Human-Control through Early Stage Intrinsic Superalignment
5 The Ultimate Superalignment: Towards Human-Superintelligence Co-alignment for Sustainable Symbiotic Society
@@ 6 Conclusion
The ethical safety risks exposed by the rapid development of AI compel us to proactively address potential superintelligence risks and establish reasonable governance and arrangements in advance to ensure AGI and Superintelligence aligning with human intentions and values. This has given rise to superalignment research, which aims to resolve the critical challenge of maintaining safety and controllability when superintelligence capabilities surpass levels of effective human supervision. Specifically, this paper conducts an in-depth analysis of the key challenges in superalignment, such as the high costs of large-scale collection of high-quality human feedback data, incomplete or failed human supervision, potential deceptive faking-alignment, and the evolving values of both humans and AI. We propose a redefined superalignment pathway that integrates external oversight and intrinsic proactive alignment.
The external oversight superalignment is built upon a dynamically adjusted, multi-level adaptive alignment approach capable of continuously aligning with humanity’s evolving value. This approach is grounded in human-centered ultimate decision, supplemented by a highly automated and interpretable value alignment evaluation system that can precisely identify misaligned scenarios, perform automatic corrections. The intrinsic proactive superalignment draws inspiration from the moral emergence mechanisms of human society, building upon the understanding of the Self and human values. By endowing Superintelligence with self-awareness and self-reflection capabilities, enabling it to distinguish good from evil, understand the harm or impact of its actions on others, incorporate empathy-based human intent inference, proactively consider human well-being, and autonomously execute ethical behavior. Achieving sustainable superalignment requires an adaptive, multi-layered ethical safeguard framework, iterative interaction and co-evolution across human values and empathy-driven AI systems, ensuring its continuous alignment with humanity’s progressively ethical and safety values. We believe that human-AI co-design and co-alignment constitute the essential pathway to developing sustainable symbiotic society.

### v4

@@ 4 Intrinsic Proactive Superalignment
Externally supervised alignment sets explicit predefined principles/supervision signals as the alignment ceiling, yet the AI’s behavior lacks genuine understanding of the constraints or the underlying essence of supervisory signals. This frequently leads to unanticipated system failures—precisely the kind of catastrophic risks that external supervision signals cannot mitigate. Thus, in addition to passive external supervision and constraints, we also need to explore internal mechanisms alignment to proactively develop and shape AI, enabling it to spontaneous align with human ethical values beyond mere external supervision.
@@ 4.2 Beneficial Meaningful Human-Control through Early Stage Intrinsic Superalignment
5 The Ultimate Superalignment: Towards Human-ASI Co-alignment for Sustainable Symbiotic Society
@@ 6 Conclusion
The ethical safety risks exposed by the rapid development of AI compel us to proactively address potential superintelligence risks and establish reasonable governance and arrangements in advance to ensure AGI and Superintelligence aligning with human intentions and values. This has given rise to superalignment research, which aims to resolve the critical challenge of maintaining safety and controllability when superintelligence capabilities surpass levels of effective human supervision. Specifically, this paper conducts an in-depth analysis of the key challenges in superalignment, such as the high costs of large-scale collection of high-quality human feedback data, incomplete or failed human supervision, potential deceptive faking-alignment, and the evolving values of both humans and AI. We propose that the value system for sustainable symbiotic society should be co-shaped and co-calibrated by both humans and AGI/ASI - an approach we define as super co-alignment, while refining the superalignment framework by integrating externally-driven oversight with intrinsically proactive alignment.
The external oversight superalignment is built upon a dynamically adjusted, multi-level adaptive alignment approach capable of continuously aligning with humanity’s evolving value. This approach is built upon a highly automated and interpretable value alignment evaluation system that can precisely identify misaligned scenarios, perform automatic corrections. The intrinsic proactive superalignment draws inspiration from the moral emergence mechanisms of human society, building upon the understanding of the Self and human values. By endowing Superintelligence with self-awareness and self-reflection capabilities, enabling it to distinguish good from evil, understand the harm or impact of its actions on others, incorporate empathy-based human intent inference, proactively consider human well-being, and autonomously execute ethical behavior. More critically, ultimate superalignment must be grounded in the co-evolution of values between humans and AGI/ASI to achieve compatibility with the value system of a sustainable symbiotic society. Should humans and ASI achieve co-alignment in co-shaping these symbiotic values, the ASI would possess sufficient rational justification for maintaining harmonious coexistence with humanity.